OpenStudy (anonymous):

What are the exact solutions of x2 − 3x − 7 = 0?

OpenStudy (inkyvoyd):

...

OpenStudy (inkyvoyd):

@friedch1cken , have you read the Code of Conduct?

OpenStudy (anonymous):

D: Sorry

OpenStudy (inkyvoyd):

Yea, you're not supposed to do that...

OpenStudy (inkyvoyd):

You can show work to the second to last step though ;)

OpenStudy (anonymous):

what about the code of conduct?

OpenStudy (inkyvoyd):

@StealerGirly50 , nothing to do with you; don't worry about it :)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[x^2 − 3x − 7 = 0\] Okay, we can begin by completing the square. Let's move the constant term to the right. Add 7 to both sides of the equation. \[x^2-3x-7+7 = 0 +7\]\[x^2-3x=7\] The x term is -3x. Take half its coefficient (1 1/2). Square it (2 1/4) and add it to both sides. Add '2 1/4' to each side of the equation. \[x^2 - 3x + 2.25= 7 +2 \frac{1}{4}\] Combine like terms. 7 + 2 1/4 = 7 1/4 \[x^2 - 3x +2 \frac{1}{4} = 9 \frac{1}{4}\] Factor a perfect square on the left side: \[(x -\frac{1}{2})(x - \frac{1}{2}) = 9 \frac{1}{4}\] \[(x-\frac{1}{2})^2 = 9 \frac{1}{4}\] We can take the square roots of both sides. \[x - \frac{1}{2} = \pm \sqrt{9 \frac{1}{4}}\] Now, we can set up two subproblems, one with a negative square root of 9.25 and the other with a positive square root of 9.25. \[x-\frac{1}{2}=\sqrt{9 \frac{1}{4}}\] Add 1/2 to both sides. \[x = \frac{1}{2} + \sqrt{9 \frac{1}{4}}\] Now we have one of our answers. We will solve the next subproblem, with a negative square root of 9/4 now. \[x- \frac{1}{2} = -\sqrt{9\frac{1}{4}}\] Add 1/2 to both sides.\[x = \frac{1}{2}-\sqrt{9\frac{1}{4}}\] And there's your second answer. So, all in all, \[x = \left\{ \frac{1}{2}-\sqrt{9\frac{1}{4}}, \frac{1}{2}+\sqrt{9 \frac{1}{4}} \right\}\] Does that make sense?

OpenStudy (anonymous):

Ya but what is supposed to be in the last (math processing error)

OpenStudy (anonymous):

What's supposed to be where?